OpenStudy (anonymous):

Solve, using the square root method: 5x2 = 60. Round the solution to the nearest tenth.

OpenStudy (saifoo.khan):

5x^2 = 60 x ^2 = 12 take square roots on both sides, \[x = \pm \sqrt{12}\]